OpenStudy (yanasidlinskiy):

What would I do to solve this? The average middle-distance runner at a local high school runs the mile in 4.5 minutes, with a standard deviation of 0.3 minute. The percentage of a runners that will run the mile in less than 4 minutes is ________%.

OpenStudy (kirbykirby):

If your data is assumed to be normally distributed, then you can find the Z score for this to help you with the problem.

OpenStudy (kirbykirby):

If it's assumed to be normal, \(X\) is the number of people that run a mile. You can standardize X by doing \[Z=\frac{X-\mu}{\sigma} \], SO the problem is asking \[ P(X < 4)\], and so y standardizing: \[P(X<4) = P\left( Z<\frac{4-4.5}{0.3}\right) =P(Z<-1.67)\]

OpenStudy (yanasidlinskiy):

I would have to solve for P?

OpenStudy (yanasidlinskiy):

Oh wait.....Is that the answer? \(\huge{\uparrow}\)